OpenStudy (anonymous):

can anybody help me for this problem 5w-6w=-w+5 please is it one solution or many solution or no solution

OpenStudy (anonymous):

No Solution

OpenStudy (lgbasallote):

no solution...coz w becomes zero when transposed

OpenStudy (anonymous):

thanks

OpenStudy (precal):

Parallel lines 5w-6w=-w+5 -w=-w+5 move +w to both sides 0=5 false statement thererfore no solution